# Grounding Multimodal LLMs to Embodied Agents that Ask for Help with Reinforcement Learning

### v2

@@ Abstract
Embodied agents operating in real-world environments must interpret ambiguous and under-specified human instructions. A capable household robot should recognize ambiguity and ask relevant clarification questions to infer the user intent accurately, leading to more effective task execution. To study this problem, we introduce the Ask-to-Act task, where an embodied agent must fetch a specific object instance given an ambiguous instruction in a home environment. The agent must strategically ask minimal, yet relevant, clarification questions to resolve ambiguity while navigating under partial observability. To solve this problem, we propose a novel approach that fine-tunes multimodal large language models (MLLMs) as vision-language-action (VLA) policies using online reinforcement learning (RL) with LLM-generated rewards. Our method eliminates the need for large-scale human demonstrations or manually engineered rewards for training such agents. We benchmark against strong zero-shot baselines, including GPT-4o, and supervised fine-tuned MLLMs, on our task. Our results demonstrate that our RL-finetuned MLLM outperforms all baselines by a significant margin ($19.1$-$40.3\%$), generalizing well to novel scenes and tasks. To the best of our knowledge, this is the first demonstration of adapting MLLMs as VLA agents that can act and ask for help using LLM-generated rewards with online RL.
Figure 1: Ask-to-Act task. In this task, the user requests a specific green cup but, instead of describing it in detail, asks the agent, “Bring the cup and place it on coffee table”. Since the user’s intent is unclear, an agent must ask a minimum number of clarification questions to disambiguate the requested object (e.g. “Are you looking for a red cup?” or “Is it on the kitchen counter?”). We consider different types of ambiguities, requiring inquiring about object attributes, spatial relationships, object size, or combinations of the three.
footnotetext: Work done as part of the internship at Meta
@@ 1 Introduction
Embodied agents operating in real-world environments must interpret human instructions, which are often ambiguous, under-specified, or context-dependent. Consider asking a household robot, ‘Bring the cup and place it on the coffee table’ To complete this task, in an environment (as shown in Fig. 1) without additional information, the robot faces multiple ambiguities like - which cup is the user referring to (red, white or green)? Does the user need a small cup or a large one? Instead of making incorrect assumptions and bringing an incorrect cup to the user, a capable household robot should ask the minimum number of clarification questions to disambiguate the requested object such as, “Are you looking for a red cup?” followed by “Is it on the kitchen counter?”. By actively asking questions that are grounded in the environment state and relevant to the context of the task, an intelligent household robot can infer user intent more accurately, leading to more effective and reliable task execution. We call this task Ask-to-Act, and believe inferring such ambiguity and asking clarification questions to resolve it requires balancing several implicit objectives such as exploration to gather task-specific information from the environment, deductive reasoning to minimize the number of questions, and contextual reasoning to determine the right time to ask questions. Humans effortlessly do this by combining common sense knowledge, contextual cues, and deductive reasoning. In this work, we ask ‘How can we build embodied agents capable of reasoning required to resolve such ambiguity by asking clarification questions?’
@@
Training such agents capable of interacting and asking clarification questions is challenging. Most tasks in embodied AI have well-defined goals [4, 46, 14, 39] and approaches for such tasks rely on either collecting large-scale human demonstrations [7, 32, 21, 27, 13] with low-level actions combined with imitation learning (IL) or use large-scale reinforcement learning [47, 2, 57] (RL) with manually designed rewards. However, neither of these methods are trivial nor scalable, especially when training agents that require intermediate natural language communication: Gathering natural language interaction data interleaved with robot actions introduces additional overhead in already expensive and time-consuming robot teleoperation. Alternatively, training such agents with RL requires access to dense reward functions, which are impractical to manually design to encourage task-specific deductive or commonsense reasoning. Prior works [33, 17, 20, 25] focusing on similar settings take advantage of the prior knowledge and rich commonsense reasoning abilities of Large Language Models (LLMs) in a zero-shot manner for task planning and reasoning about ambiguity to ask clarification questions. However, these methods either require careful prompt engineering to prevent the robot from excessively relying on seeking assistance, or operate under strong assumptions, i.e. the environment is fully observable and can be represented in text without any errors. These assumptions are unrealistic – real-world environments impose partial observability, requiring agents to actively explore and gather information.
@@
To address these limitations, our key idea is to bootstrap the learning signal required for training embodied agents capable of interacting and asking in an end-to-end manner by leveraging an LLM’s contextual commonsense and deductive reasoning ability to resolve ambiguity. We accomplish this by adapting a multimodal large language model [22, 23] (MLLM) into a vision-language-action [43, 7, 21, 41] (VLA) model using large-scale RL with reward signal generated using an LLM with access to privileged information from the simulator. To distill reasoning ability required for resolving ambiguity into this VLA model, we propose to use per-step rewards generated using an LLM [24, 54] that evaluates both the actions taken and natural language question asked by the agent in context of the task. Through our experiments, we show that LLMs are highly effective at generating per-step rewards for such tasks that require interacting with environment and asking questions to resolve ambiguity, when provided with the right representation of task and environment in text – information that can be easily curated at training time using privileged simulator state. This framework enables us to adapt MLLMs into VLA models that can interact with the environment and resolve ambiguity by asking questions without expensive human demonstrations or manually engineered rewards.
@@
To evaluate effectiveness of our method we instantiate the Ask-to-Act task in a more realistic setup where agents operate under partial observability, process visual observations, and at each step output either actions or natural language questions to solve a rearrangement task [46, 39], in contrast to prior works [33, 20]. We set up Ask-to-Act in Habitat 3.0 [30] using 83 scenes from the ReplicaCAD [39] dataset, with 42 object categories from Google Scanned Objects [12]. We evaluate the generalization ability of these agents along two axes: 1) Unseen Scenes: where we evaluate on seen tasks in unseen scenes which include unseen configurations of furniture and objects, and 2) Unseen Tasks: where we evaluate the agent on novel task instances in unseen scenes i.e. novel compositions of the five types of ambiguity described in Sec. 3 not seen by the agent during training.
@@
We compare our proposed method against strong zero-shot baselines leveraging proprietary LLMs and MLLMs like GPT-4o [26], as well as methods for fine-tuning open-source MLLMs using synthetically generated supervised fine-tuning (SFT) data for the Ask-to-Act task. Our method, which employs large-scale RL with LLM-generated rewards to fine-tune an open-source MLLM, outperforms all baselines by a large margin, $40.3\%$ and $19.1\%$ for success rates on Unseen Scenes and Unseen Tasks evaluation splits. We also find that training VLA policies using only task subgoal rewards (i.e. no explicit rewards for asking the right questions), which can be programmatically written, achieve performance close to random on Ask-to-Act task. This indicates that training embodied agents to act and reason about ambiguity requires dense reward signals. Additionally, we analyze the behavior of our RL-trained policy and find that as the number of questions these agents can ask increases, the success rates of these policies improve and they generalize better to Unseen Tasks.
@@ 2 Related Work
Agents that ask for help. Our work builds on prior works in embodied AI [28, 45, 15, 3, 38] which explore effective human-robot interaction through dialogue. In these settings, an embodied agent can ask questions in natural language to improve task performance or human-robot interaction. In Ask-to-Act, we study this in the context of rearrangement tasks specified with under-specified instructions where the agent needs to reason about object appearance, location, and size, engaging in a multi-round dialogue (1-4 rounds) to disambiguate the object. Similar to prior works [28, 45, 15], we restrict the vocabulary of questions in order to make it feasible to evaluate the questions agent asks. In recent works [33, 20], the most common approach is to leverage an LLM’s commonsense reasoning and language generation ability in a zero-shot manner to handle such human-robot interactions. However, these methods require extensive prompting and operate under assumption of full observability i.e. the complete error-free environment state is represented in text. We take a step towards relaxing these constraints and make the task setup more realistic by setting up the Ask-to-Act task in a partially observable setting.
Vision-Language-Action Models. Prior works [43, 41, 42, 21, 8, 7] have demonstrated that multimodal large language models can be adapted as vision-language-action (VLA) models for interactive decision-making tasks. These methods leverage MLLMs [6, 22, 23] pre-trained on vision-language modeling tasks and finetune them for low-level robot control [21, 7, 42, 41] or embodied task planning [43, 48, 56]. These works focus on settings where a task is clearly specified and there is no ambiguity which makes evaluation and training data/signal generation easier. In our work, we focus on following the same recipe of finetuning MLLMs for embodied task planning but for tasks that requires contextual reasoning and asking for external feedback from a user for information not accessible from observing the environment.
Reward generation using LLMs. A promising method to train robot policies without expensive human demonstrations is to use RL with synthetically generated rewards [24, 55, 9, 34, 24]. A common approach is to prompt the LLM to write code that takes symbolic features from the environment observations and produces a scalar output representing the reward. This approach has shown promising results on training policies for low-level robot control for complex dexterous manipulation, whole body control, locomotion, and object manipulation. A key difference of our work is that we are focused on teaching embodied agents ambiguity reasoning ability and language generation to ask for feedback using LLM generated rewards. Motivated by the effectiveness of using LLMs as reward models for complex math and code reasoning tasks [58, 34, 49, 10], in our work we focus on evaluating the effectiveness of LLM reward models for embodied tasks that require reasoning about ambiguity. In our work, we show that this paradigm of training interactive policies using an LLM as a process reward model (PRM) [37] is an effective approach for tasks that require reasoning about ambiguity, where writing rewards manually or collecting human demonstrations is not trivial or scalable.
@@ 3 Ask-to-Act task
Question Query Type
Is it the [Attr] [Obj]? Object attribute
Is the object on the [Recep]? Object location
Is it the [Size] [Obj]? Object size
Table 1: Types of questions agent can ask in Ask-to-Act task.
@@
Our goal is to build embodied agents capable of reasoning about ambiguity when task-planning and interacting with users for clarification when required under partially observable settings. To study building such agents, we propose the Ask-to-Act task as shown in Fig. 1, where an agent is spawned randomly in an unseen indoor environment and tasked with fetching a specific instance of an object and placing it on a designated receptacle using an underspecified language instruction. Consider the example shown in Fig. 1: the user wants a specific green cup but, instead of describing it in detail, asks the agent, “Bring the cup and place it on coffee table”. While searching, the agent finds multiple cups on the kitchen counter and the table - two red cups, a white cup, and two green cups. Since the user’s intent is unclear, an optimal agent must ask the fewest clarification questions to disambiguate the requested object. To study an agent’s ability to reason about ambiguity, we synthetically create tasks with different types of ambiguities, including variations in object appearance, location, and size. As shown in Fig. 1, our task evaluates ambiguity across five dimensions:
@@
  • Attribute Recognition: The agent must reason about appearance-based attributes such as color and object category to disambiguate objects (e.g., a blue shoe, white shoe and orange shoe as shown in Fig. 1 (a)).
@@
  • Spatial Reasoning: The agent must differentiate between objects with similar appearance and geometry based on their spatial location (e.g., red cup on the light table, or kitchen counter, or dark table as shown in Fig. 1 (b)).
@@
  • Object Size: The agent must distinguish between objects of different sizes but similar appearance and location (e.g., a large red bowl vs. a small red bowl).
@@
  • Attribute and Spatial Reasoning: A combination of attribute recognition and spatial reasoning, where the agent must consider both appearance and location to minimize the number of clarification questions needed to identify the target object (as shown by example in Fig. 1 (c)).
  • Attribute/Spatial Reasoning and Object Size: A combination of attribute and spatial reasoning with object size, where the agent must account for both appearance or location and size to effectively disambiguate the target object.
These five axes of ambiguity comprehensively assess various capabilities, including ones that MLLMs are known to struggle with (e.g. counting [11], and spatial reasoning [31]).
Evaluation Splits. We assess the generalization ability of these agents along two axes:
  • Unseen Scenes: Evaluates how well the agent can perform tasks under new object layouts in the scene, or new instances of previously seen object classes.
@@
  • Unseen Tasks: Tests the agent’s ability to handle novel instances of ambiguity in unseen scenes, requiring the agent to ask more questions than training tasks. For example, for attribute recognition tasks, in training the agent only needs to distinguish between 2 objects of different attributes but in Unseen Tasks split it needs to distinguish between $3$ or $4$ instances. See Appendix A for examples.
Evaluating Questions. Evaluating questions asked by an agent in context of a task in a programmatic manner is important for simulating the user response. To do so, we use an LLM to build an answering module for Ask-to-Act task which takes as input the task instruction, privileged information of state of environment represented as text, metadata of the target object a user is requesting, and is tasked with answering questions generated by the embodied agent. In order to constrain the difficulty of natural language question generation and answering, we limit the types of questions an agent can ask to $3$ types of questions (via prompting) that query object attribute, location, and size as shown in Tab. 1. We apply this constraint to ensure we can accurately evaluate questions asked by an agent as open-ended question-answering is an open research problem. For each question an agent asks, the LLM responds with either a ‘Yes/No’ language response. When training MLLM policies using RL, having an LLM as answering module and reward model in the training loop introduces additional infrastructure challenges and increases training time dramatically. By constraining the types of questions, we can compute the answers and rewards from an LLM for our task offline which allows to scale RL training. Details about the implementation are in Appendix D.
@@
Task Setup. We instantiate the task using Habitat 3.0 [30] in $63$ training scenes and $20$ evaluation scenes from ReplicaCAD [39] dataset with $42$ object categories from Google Scanned Objects [12]. We use the Spot robot embodiment for the Ask-to-Act task with a height of $1.41$m and radius of $0.25$cm. At each timestep the agent has access to a $480\times 640$ resolution RGB image, robot joint positions, an indicator of whether the robot is holding a object or not, relative base egomotion, and task instruction.
Action Space. The agent’s action space is implemented in Habitat 3.0 [30]. We provide all methods with a set of oracle low-level skills, and focus on high-level decision-making. We opt for a hierarchical approach for our task, as training end-to-end with low-level actions for such a task presents the challenge of long-horizon task execution that requires tens of thousands steps of interaction while also engaging in natural language interaction with the user. Even seemingly simple mobile pick-and-place tasks remain difficult for end-to-end approaches [5, 18]. However, hierarchical methods - where a high-level policy selects from a set of predefined skills have proven effective in rearrangement tasks [40, 39, 1]. At each timestep, the high-level policy selects from 116 predefined skills or an open-ended “ask question” action, to generate a natural language query when clarification is needed. These skills include picking objects by name, placing on receptacles, and navigating to specific locations. Example action outputs include pick(apple), place(sink), navigate(coffee table) and ask_question(“Do you want the red cup?”).

### v3

@@ Abstract
Embodied agents operating in household environments must interpret ambiguous and under-specified human instructions. A capable household robot should recognize ambiguity and ask relevant clarification questions to infer the user intent accurately, leading to more effective task execution. To study this problem, we introduce the Ask-to-Act task, where an embodied agent is tasked with a single or multi-object rearrangement task using an under-specified instruction in a home environment. The agent must strategically ask minimal, yet relevant, clarification questions to resolve ambiguity while navigating under partial observability. To address this challenge, we propose a novel approach that fine-tunes multi-modal large language models (MLLMs) as vision-language-action (VLA) policies using online reinforcement learning (RL) with LLM-generated rewards. Our method eliminates the need for large-scale human demonstrations or manually engineered rewards for training such agents. We benchmark against strong zero-shot baselines including GPT-4o as well as supervised fine-tuned MLLMs on our task. Our results show that our RL-finetuned MLLM outperforms all baselines by a significant margin ($10.4$-$16.5\%$), generalizing well to novel scenes and tasks. To the best of our knowledge, this is the first demonstration of adapting MLLMs as VLA agents that can act and ask for help using LLM-generated rewards with online RL.
@@ 1 Introduction
Embodied agents interacting with humans must interpret human instructions that are often ambiguous, under-specified, or context-dependent. Consider asking a household robot, ‘Bring the cup and place it on the coffee table’ To complete this task in an environment (as shown in Fig.˜1) without additional information, the robot faces multiple ambiguities like - which cup is the user referring to (red, white or green)? Does the user need a small cup or a large one? Instead of making incorrect assumptions and bringing an incorrect cup to the user, a capable household robot should ask the minimum number of clarification questions to disambiguate the requested object such as, “Are you looking for a red cup?” followed by “Is it on the kitchen counter?”. By actively asking questions that are grounded in the environment state and relevant to the context of the task, an intelligent household robot can infer user intent and preferences more accurately, leading to more effective and reliable task execution. We call this task Ask-to-Act, and believe inferring such ambiguity or user preference and asking clarification questions to resolve it requires balancing several implicit objectives such as exploration to gather task-specific information from the environment, deductive reasoning to minimize the number of questions, and contextual reasoning to determine the right time to ask questions. Humans effortlessly do this by combining common sense knowledge, contextual cues, and deductive reasoning. In this work, we ask ‘How can we train embodied agents capable of reasoning required to resolve ambiguity and user preferences by asking clarification questions?’
Figure 1: Ask-to-Act task. In this task, the user requests a specific green cup but, instead of describing it in detail, asks the agent, “Bring the cup and place it on coffee table”. Since the user’s intent is unclear, an agent must ask a minimum number of clarification questions to disambiguate the requested object (e.g. “Are you looking for a red cup?” or “Is it on the kitchen counter?”). We consider under-specified single and multi-object rearrangement tasks that involve inquiring about user preferences and resolving different types of ambiguities, about object attributes, spatial relationships, object size, placement location, or combinations of the four.
@@
Training such agents capable of interacting and asking clarification questions is challenging. Most tasks in embodied AI have well-defined goals (Batra et al., 2020; Weihs et al., 2021; Ehsani et al., 2021; Szot et al., 2021) and approaches for such tasks rely on either collecting large-scale human demonstrations (Black et al., 2024; Ramrakhya et al., 2022; Kim et al., 2024; O’Neill et al., 2024; Ehsani et al., 2024) with low-level actions combined with imitation learning (IL) or use large-scale reinforcement learning (Wijmans et al., 2020; Andrychowicz et al., 2018; Zeng et al., 2024) (RL) with manually designed rewards. However, neither of these methods are trivial nor scalable, especially when training agents that require intermediate natural language communication: Gathering natural language interaction data interleaved with robot actions introduces additional overhead in already expensive and time-consuming robot teleoperation. Alternatively, training such agents with RL requires access to dense reward functions, which are impractical to manually design to encourage task-specific deductive or commonsense reasoning. Prior works (Ren et al., 2023; Huang et al., 2022; Jr. & Manocha, 2024; Min et al., 2024) focusing on similar settings take advantage of the prior knowledge and rich commonsense reasoning abilities of Large Language Models (LLMs) in a zero-shot manner for task planning and reasoning about ambiguity to ask clarification questions. However, these methods either require careful prompt engineering to prevent the robot from excessively relying on seeking assistance, or operate under strong assumptions, i.e. the environment is fully observable and can be represented in text without any errors. These assumptions are unrealistic – real-world environments impose partial observability, requiring agents to actively explore and gather information.
@@
To address these limitations, our key idea is to bootstrap the learning signal required for training embodied agents capable of interacting and asking in an end-to-end manner by leveraging an LLM’s contextual commonsense and deductive reasoning ability to resolve ambiguity. We accomplish this by adapting a multimodal large language model (MLLM) (Li et al., 2024; Liu et al., 2024) into a vision-language-action (VLA) model (Szot et al., 2024b; Black et al., 2024; Kim et al., 2024; Szot et al., 2024a) using large-scale RL with reward signal generated using an LLM. To distill reasoning ability required for resolving ambiguity into this VLA model, we propose to use per-step rewards generated using an LLM (Ma et al., 2024; Yu et al., 2023a) that evaluates both the actions taken and natural language question asked by the agent in context of the task. Through our experiments, we show that LLMs are highly effective at generating per-step rewards for such tasks that require interacting with environments and asking questions, when provided with the right representation of task and environment in text – information that can be easily curated at training time using privileged simulator state. Unlike prior works (Ma et al., 2024; Yu et al., 2023a) that use LLMs to generate code based rewards for low-level control policies, our work is one of the first ones to demonstrate LLMs can be used as reward model for tasks that require common sense and deductive reasoning in embodied settings.
@@
To evaluate the effectiveness of our method we instantiate the Ask-to-Act task in a partial observable setup where agents process visual observations, and at each step output either actions or natural language questions to solve a rearrangement task (Weihs et al., 2021; Szot et al., 2021), in contrast to prior works (Ren et al., 2023; Jr. & Manocha, 2024). We set up Ask-to-Act in Habitat 3.0 (Puig et al., 2024) using 83 scenes from the ReplicaCAD Szot et al. (2021) dataset, with 42 object categories from Google Scanned Objects (Downs et al., 2022). We evaluate the generalization ability of these agents along two axes: 1) Unseen Scenes: where we evaluate on seen tasks in unseen scenes which include unseen configurations of furniture and objects, and 2) Unseen Tasks: where we evaluate the agent on novel task instances in unseen scenes i.e. novel compositions of the tasks with ambiguity described in Sec.˜3 not seen by the agent during training.
@@
We compare our proposed method against strong zero-shot baselines leveraging proprietary LLMs and MLLMs like GPT-4o (OpenAI, 2023), as well as methods for fine-tuning open-source MLLMs using synthetically generated supervised fine-tuning (SFT) data for the Ask-to-Act task. Our method, which employs large-scale RL with LLM-generated rewards to fine-tune an open-source MLLM, outperforms all baselines by a large margin, $10.4\%$ and $16.5\%$ for success rates on Unseen Scenes and Unseen Tasks evaluation splits. We also find that training VLA policies using only task subgoal rewards (i.e. no explicit rewards for asking the right questions), which can be programmatically written, achieve performance close to random on Ask-to-Act task. This indicates that training embodied agents to act and reason about ambiguity requires dense reward signals. Additionally, we analyze the behavior of our RL-trained policy and find that as the number of questions these agents can ask increases, the success rates of these policies improve and they generalize better to Unseen Tasks.
@@ 2 Related Work
Agents that ask for help. Our work builds on prior works in embodied AI (Padmakumar et al., 2021; Thomason et al., 2019; Gao et al., 2022; Banerjee et al., 2020; Singh et al., 2022) which explore effective human-robot interaction through dialogue. In these settings, an embodied agent can ask questions in natural language to improve task performance. In Ask-to-Act, we study this in the context of multi-object rearrangement tasks with under-specified instructions where the agent needs to reason about ambiguity and identify user preferences by engaging in a multi-round dialogue (1-7 rounds). In recent works (Ren et al., 2023; Jr. & Manocha, 2024), the most common approach is to leverage an LLM’s commonsense reasoning and language generation ability in a zero-shot manner to handle such human-robot interactions. However, these methods require extensive prompting and operate under assumption of full observability i.e. the complete error-free environment state is represented in text. We take a step towards relaxing these constraints and make the task setup more realistic by setting up the Ask-to-Act task in a partially observable setting.
Reward generation using LLMs. A promising method to train robot policies without expensive human demonstrations is to use RL with synthetically generated rewards (Ma et al., 2024; Yu et al., 2023b; Chen et al., 2025; Sarukkai et al., 2024; Ma et al., 2024). A common approach is to prompt the LLM to write code that takes symbolic features from the environment observations and produces a scalar output representing the reward. This approach has shown promising results on training policies for low-level robot control for complex dexterous manipulation, whole body control, locomotion, and object manipulation. A key difference of our work is that we are focused on teaching embodied agents ambiguity reasoning ability and language generation to ask for feedback using LLM generated rewards. Motivated by the effectiveness of using LLMs as reward models for complex math and code reasoning tasks (Zeng & Xu, 2023; Sarukkai et al., 2024; Xie et al., 2024; Chu et al., 2023), in our work we focus on evaluating the effectiveness of LLM reward models for embodied tasks that require reasoning about ambiguity. In our work, we show that this paradigm of training interactive policies using an LLM as a process reward model (PRM) (Setlur et al., 2025) is an effective approach for tasks that require reasoning about ambiguity, where writing rewards manually or collecting human demonstrations is not trivial or scalable.
@@ 3 Ask-to-Act task
Our goal is to build embodied agents capable of completing under-specified tasks by interacting with users to ask clarification questions to resolve ambiguity and identify preferences of a user in partially observable settings. To study building such agents, we propose the Ask-to-Act task as shown in Fig.˜1, where an agent is spawned randomly in an unseen indoor environment and tasked with fetching single or multiple instances of objects and placing them on a designated receptacle using an under-specified language instruction. Consider the example shown in Fig.˜1: the user wants a specific green cup but, instead of describing it in detail, asks the agent, “Bring the cup and place it on coffee table”. While searching, the agent finds multiple cups on the kitchen counter and the table - two red cups, a white cup, and two green cups. Since the user’s intent is unclear, an optimal agent must ask the fewest clarification questions to disambiguate the requested object. To study an agent’s ability to reason about ambiguity and preferences of the user for such under-specified tasks, we synthetically create tasks of varying difficulty that require identifying user preferences or resolving different types of ambiguities, e.g. about variations in object appearance, location, placement location, size or composition of these axes. Fig.˜1 and Fig.˜5 shows examples of tasks in our dataset, and below we describe each task and evaluation criteria in detail:
1.) No Ambiguity: These include single or multi-object rearrangement instructions that clearly specify the target objects and do not require reasoning about ambiguity or user preferences.
@@
2.) Attribute Recognition: Agent must reason about appearance attributes like color or category to disambiguate target objects (e.g., blue shoe, white shoe vs. orange shoe as shown in Fig.˜1 (a)).
@@
3.) Spatial Reasoning: Agent must differentiate between objects with similar appearance based on spatial location to identify target object (e.g., red cup on the light table, or kitchen counter, or dark table as shown in Fig.˜5 (b) in Appendix˜A).
@@
4.) Object Size: The agent must distinguish between objects of different sizes but similar appearance and location to identify target object (e.g., a large red bowl vs. a small red bowl, shown in Fig.˜5 (e) in Appendix˜A).
@@
5.) Compositional Ambiguity Reasoning: A combination of attribute recognition, object size, and spatial reasoning, where the agent must consider appearance, geometry and location to ask minimum number of clarification questions to identify the target object (as shown by example in Fig.˜5 (c, f) in Appendix˜A).
6.) Clean Clutter: Multi-object rearrangement tasks with under-specified instructions, where the agent must determine which objects on a source receptacle are “clutter” and place them in appropriate target locations. In each instance, the user does not know the full set of objects on the source receptacle, and only a subset of the objects present are considered clutter. As a result, the agent must communicate the current state of the receptacle (see example in Fig.˜1 (b)). Additionally, the user has hidden placement preferences for each object category (e.g., cups should go in the left drawer of the kitchen counter, while plates belong in the top cabinet).
7.) Preference based: Multi-object rearrangement tasks that require identifying user preferences of placement location for multiple objects of a semantic category. For example, “Move all utensils to kitchen” requires asking user for preference of target receptacle for each object categorized as utensils (e.g. cups go on top drawer and bowls go on right drawer of kitchen counter, as shown in Fig.˜1 (c)).
Simulation Answers. Evaluating questions asked by an agent in context of a task in Ask-to-Act in a programmatic manner is important for simulating diverse users preferences and language responses. To do so, we use an LLM to build an answering module for Ask-to-Act task which takes as input the task instruction, privileged information of environment state represented as text, metadata of the target objects a user is requesting and expected target location, and is tasked with answering questions generated by the embodied agent. Details about the prompts to answering LLM are in Appendix˜C
@@
Evaluation Splits. We assess the generalization ability of these agents along two axes: (1.) Unseen Scenes: Evaluates how well the agent can perform tasks under new object layouts in the scene, or new instances of previously seen object classes. (2.) Unseen Tasks: Tests the agent’s ability to handle novel instances of ambiguity in unseen scenes, requiring the agent to ask more questions than training tasks. For example, for attribute recognition tasks, in training the agent only needs to distinguish between 2 objects of different attributes but in Unseen Tasks split it needs to distinguish between $3$ or $4$ instances. See Appendix˜A for examples.
@@
Task Setup and Agent Action Space. We instantiate the task using Habitat 3.0 (Puig et al., 2024) in $63$ training scenes and $20$ evaluation scenes from the ReplicaCAD (Szot et al., 2021) dataset with $42$ object categories from Google Scanned Objects (Downs et al., 2022). We use the Spot robot embodiment for the Ask-to-Act task with a height of 1.411.41m and radius of 0.250.25cm. At each timestep the agent has access to a $480\times 640$ resolution RGB image, robot joint positions, an indicator of whether the robot is holding an object or not, relative base egomotion, and task instruction. The agent’s action space is implemented in Habitat 3.0 (Puig et al., 2024) with a set of oracle low-level skills, and focus on high-level decision-making. Low-level skills include picking objects by name, placing on receptacles, navigating to specific locations, and opening articulated receptacles. Example action outputs include pick(apple), place(sink), navigate(cabinet), open(left drawer), and ask_question(“Do you want the red cup?”). At each timestep, the policy selects from 125 predefined skills or an open-ended “ask question” action to generate a natural language query when clarification is needed.